# ROSplane 2.0: A Fixed-Wing Autopilot for Research

### v1

@@ Abstract
Unmanned aerial vehicle (UAV) research requires the integration of cutting-edge technology into existing autopilot frameworks. This process can be arduous, requiring extensive resources, time, and detailed knowledge of the existing system. ROSplane is a lean, open-source fixed-wing autonomy stack built by researchers for researchers. It is designed to accelerate research by providing clearly defined interfaces with an easily modifiable framework. Powered by ROS 2, ROSplane allows for rapid integration of low or high-level control, path planning, or estimation algorithms. A focus on lean, easily understood code and extensive documentation lowers the barrier to entry for researchers. Recent developments to ROSplane improve its capacity to accelerate UAV research, including the transition from ROS 1 to ROS 2, enhanced estimation and control algorithms, increased modularity, and an improved aerodynamic modeling pipeline. This aerodynamic modeling pipeline significantly reduces the effort of transitioning from simulation to real-world testing without requiring expensive system identification or computational fluid dynamics tools. ROSplane’s architecture reduces the effort required to integrate new research tools and methods, expediting hardware experimentation.
@@ I Introduction
Researchers often face significant integration problems when it comes to conducting realistic simulations and real-world flight tests. Research software can require extensive rewriting or refactoring to be compatible with realistic simulations or hardware platforms. This increases the difficulty of conducting real-world experiments, which are essential steps in validating and proving out new research algorithms and methods.
ROSflight[1] is a lean, open-source [^1] https://github.com/rosflight autopilot designed to mitigate these challenges and reduce the barrier to entry for UAV research. It accomplishes this by allowing the same software that runs in simulation to also control the physical vehicle with no changes. Since ROSflight is built on the robot operating system (ROS 2), it offers superior modularity and customizability.
@@ II Related Work
Many excellent autopilots and flight control software are available for hobby, commercial, and research users alike. Current open-source autopilots like PX4[4] and ArduPilot [5] offer excellent plug-and-play capability. They have the advantage of years of development, large communities, and full-featured autonomy stacks. Although they offer distinct advantages, these autopilots have large code bases not designed for easy modification or integration, and can require a researcher to gain familiarity with these large code bases to implement given research. It can also lead to a black box environment by obfuscating core autopilot features, making integration of research code difficult and time consuming. Furthermore, code for these autopilots runs primarily on embedded microcontrollers, which makes debugging and active development more difficult and microcontroller dependent.
@@
ROSplane 2.0 offers an autonomy stack that is designed for easy integration with fixed-wing UAV research of any kind. Its lean feature set allows the code to be easily understood, which lowers the integration time and effort for researchers, thus also lowering the barrier to entry and the learning curve. ROSplane also moves the entirety of the autopilot stack from the embedded microcontroller-based flight control unit (FCU) to the Linux-based companion computer. This facilitates research code development and means that no change of firmware-level code is required, the code is easily modifiable, and transition from simulation to flight is seamless. Pairing this with powerful tools like Docker [7] makes testing and development transitions from vehicle to vehicle effortless. The ROSplane architecture is described more fully in Section III.
The Robot Operating System (ROS 2) has been extensively used for research and commercial use alike[8]. Existing autopilots like PX4 have made significant efforts to support ROS 2 software, and integration with ROS 2 software and PX4 is constantly improving [6]. ROSplane takes ROS 2 support a step further by structuring all autopilot features as ROS 2 nodes, including controllers, estimators, and path planners. This tight integration with ROS 2 significantly improves customizability and modularity.
@@ III-B Autopilot Structure
This architecture gives ROSplane basic waypoint-following functionality that users can use out-of-the-box. Due to the design philosophy of ROSplane, the waypoint-following functionality is neither complex, advanced, nor fully-featured. Instead, the structure of ROSplane facilitates easy understanding with clearly delineated responsibilities assigned to each module and lean, well-understood implementations. This makes it easier for users to identify which modules are necessary for their specific applications and which modules should be removed, replaced, or modified, thus increasing modularity and customizability while reducing time invested in deployment.
@@ V-A State Estimation
Previously, ROSplane used a pedagogically motivated two-stage estimation scheme. It separated attitude estimation and positional states. This made it easier for students to implement [2]; however, the estimator was less accurate than desired. In ROSplane 2.0, the estimator has been replaced with a full-state extended Kalman filter (EKF). The EKF has also expanded the number of states being estimated, including heading, gyroscope biases and lateral velocity. The estimator also takes full advantage of GNSS velocity measurements and measurements from the magnetometer. Figure 3 compares the ROSplane and ROSplane 2.0’s estimators on simulated data. The EKF has significantly less tracking error in attitude, velocity and position, and demonstrates that it provides consistent and reliable estimation to support the flight experiments of researchers.
Figure 3: Comparison of ROSplane and ROSplane 2.0 estimated states on a takeoff and path following sequence in simulation.
The EKF has also had modularity improvements. It is designed to easily accommodate new measurement updates, allowing the researcher to easily integrate new sensors. The framework for adding new sensors can be found on the project website.
@@ V-B Control
The controller’s modularity improvements allow for safe integration of new control schemes. This includes utilizing a state machine to ensure that full autonomous control is only taken at safe altitudes. ROSplane 2.0 has a simple state machine to allow for the controller to have different control schemes based on the altitude and phase of flight. These are summarized in Figure 4. Integration with these phases of flight and the modular nature of the controller maximize flexibility for the researcher to integrate research by allowing them to take as much or as little control as they need.
@@ VI-C Intended Workflow for ROSplane
This section describes the intended workflow for users interested in ROSplane. The first step is for a researcher to develop their application-specific code and integrate it into the ROSplane autonomy stack. Then, the aircraft used in hardware flight tests should be modeled as described in Section IV. If hardware tests will not be performed, then the default aerodynamic model can be used. It is recommended that users then use the ROSflight simulation to test and further refine the application code. The project website has detailed information on setting up the simulation environment. Once the application code works well in simulation, hardware tests can be performed. Because ROSflight and ROSplane run the exact same way in simulation and in hardware [1], these hardware tests can be performed in the exact same manner as the simulation tests were performed.
@@ VII Results
This section presents the results of hardware and simulation experiments demonstrating the basic functionality ROSplane offers to users out-of-the-box. These results also demonstrate that ROSplane closes the simulation-to-real gap for research. In each test, the aircraft was directed to fly to a series of four waypoints in an hourglass formation on a continuous circuit.
VII-A Hardware
To validate the performance of ROSplane on standard UAV hardware, several maneuvers were flown that tested the performance of the controller and estimator. The desired path and the performance can be seen in Figure 5. The path following algorithm overshoots the turns of the path because the commanded path assumes that the aircraft is only acceleration-limited and can achieve its maximum bank angle instantaneously. Figure 6 shows that the aircraft was able to respond to given commands effectively. Times of poor response in the pitch loop are seen to be coincident to periods of high roll command. It is important to note that the oscillations seen in the roll response have a frequency on the order of four seconds and are in response to deviations of the course from the commanded and are not due to instability in the roll controller. The amplitude of the roll response oscillations is small, typically less than 15 degrees. Similarly, the course oscillations have amplitudes less than 10 degrees.
The altitude tracking often dips well below and above the actual target. This is because as the aircraft banks, the effective lift is decreased, and the pitch control loop takes time to compensate. The aircraft overshoots its target altitude as well. This error is likely contributed to by pressure effects from the barometer measuring cabin pressure rather than true atmospheric, since the FCU is housed inside the fuselage of the aircraft. This is manifest as a decreasing error between the RTK and estimated down position as the aircraft decreases in altitude.
The estimator accurately estimates the position of the aircraft and yields an RMSE from RTK in the north east position estimation of 1.93 meters. The total RMSE error for the estimated path is 3.64 meters. Though no ground truth was available for the other states, the estimator gives reasonable and expected outputs. The estimation of these states can be seen in Figures 5 and 6.
Figure 5: Comparison of ROSplane estimated position and RTK GNSS recorded position, alongside commanded path.
Figure 6: The controller response to given flight path input in the airspeed and attitude angles.
These results show that ROSplane is an effective platform for research. Researchers can implement desired algorithms easily to evaluate different control, estimation, trajectory planning and following algorithms. Full discussion and analysis of the controller and rationale for its performance can be found in [3].
@@ VII-B Simulation to Real Flight
The simulation capabilities of ROSplane and the aerodynamic modeling pipeline were tested according to the aircraft integration process outlined in Section IV. The responses of the controller in simulation were then compared to those recorded during physical flight, given the same trajectory commands.
@@
A model of an RMRC Anaconda UAV aircraft was created in both XFLR5 and OpenVSP, and a variety of aerodynamic and stability analyses were performed to obtain the aerodynamic coefficients and stability and control derivatives for the aircraft. XFLR5 and OpenVSP both produced aerodynamic models that were accurate enough for controller tuning, though each tool excelled in different areas. XFLR5’s built-in trimming capability made stability analyses straightforward and yielded more consistent stability derivatives. OpenVSP, by modeling the full aircraft geometry, provided more reliable drag estimates and control derivatives. For both the Xflr5 and OpenVSP models, several cycles of iteration, as outlined in Section IV, were required for the aircraft model to converge on realistic performance. The final performance of the simulation and physical flight was comparable. Note that the same controller gains were used in both flights, meaning that there was no difference in ROSplane’s controllers for this comparison. Figure 8 shows the OpenVSP simulation results compared with several physical flight test results. The Xflr5 results were similar.
@@
These comparisons validate that new UAVs can be modeled and analyzed using open-source software, integrated into the ROSplane simulation, and used for controller tuning, mission planning, and research testing. Iterative testing and improvement of these models provided confidence that real-life system performance can be accurately predicted within a safe simulation environment.
@@ VIII Conclusion
ROSplane’s structure, capabilities, and modularity provide researchers with more flexibility and lower the barriers to entry to perform UAV autonomy research. The improved code structure and architecture provide greater customizability and opportunities for integrating external hardware and code. ROSplane 2.0 architecture allows researchers to quickly and safely tune flight controllers for new fixed-wing UAVs in simulation to prepare for physical flight tests. This functionality of ROSplane 2.0 was demonstrated by modeling a UAV aircraft using open-source aircraft modeling and analysis software, integrating this aircraft model into the ROSplane simulation, and validating that the closed-loop controller performance of the simulation and physical flight are comparable. This enables researchers to perform accurate controller tuning in simulation before flight, thereby reducing risk and iteration time. The universal applicability of the ROS 2 structure, paired with proven functionality and an effective simulation platform, makes ROSplane a powerful tool for researchers valuing customizability and direct control in cutting-edge autonomous research applications.

### v2

@@ Abstract
Unmanned aerial vehicle (UAV) research requires the integration of cutting-edge technology into existing autopilot frameworks. This process can be arduous, requiring extensive resources, time, and detailed knowledge of the existing system. ROSplane is a lean, open-source fixed-wing autonomy stack built by researchers for researchers. It is designed to accelerate research by providing clearly defined interfaces with an easily modifiable framework. Built around ROS 2, ROSplane allows for rapid integration of low or high-level control, path planning, or estimation algorithms. A focus on lean, easily-understood code and extensive documentation lowers the barrier to entry for researchers. Recent developments to ROSplane improve its capacity to accelerate UAV research, including the transition from ROS 1 to ROS 2, enhanced estimation and control algorithms, increased modularity, and an improved aerodynamic modeling pipeline. This aerodynamic modeling pipeline significantly reduces the effort of transitioning from simulation to real-world testing without requiring costly system identification or computational fluid dynamics tools. ROSplane’s architecture reduces the effort required to integrate new research tools and methods, expediting hardware experimentation.
@@ I Introduction
Researchers often face significant integration challenges when it comes to conducting realistic simulations and real-world flight tests. Research software can require extensive rewriting or refactoring to be compatible with realistic simulations or hardware platforms. This increases the difficulty of conducting real-world experiments, which are essential steps in validating and proving out new research algorithms and methods.
ROSflight[11] is a lean, open-source [^1] https://github.com/rosflight autopilot designed to mitigate these challenges and reduce the barrier to entry for UAV research. It accomplishes this by allowing the same software that runs in simulation to also control the physical vehicle with no changes. Since ROSflight is built on the Robot Operating System (ROS 2)[7], it offers superior modularity and customizability.
@@ II Related Work
Many excellent autopilots and flight control software are available for hobby, commercial, and research users alike. Current open-source autopilots like PX4[8] and ArduPilot [2] offer excellent plug-and-play capability. They have the advantage of years of development, large communities, and full-featured autonomy stacks. Although they offer distinct advantages, these autopilots have large code bases not designed for easy modification or integration, and can require a researcher to gain familiarity with these large code bases to implement given research. It can also lead to a black box environment by obfuscating core autopilot features, making integration of research code difficult and time consuming. Furthermore, code for these autopilots runs primarily on embedded microcontrollers, which makes debugging and active development more difficult and microcontroller-dependent.
@@
ROSplane 2.0 offers an autonomy stack that is designed for easy integration with fixed-wing UAV research of any kind. Its lean feature set allows the code to be easily understood, which lowers the integration time and effort for researchers, thus also lowering the barrier to entry and the learning curve. ROSplane also moves the entirety of the autopilot stack from the embedded microcontroller-based flight control unit (FCU) to the Linux-based companion computer. This facilitates research code development and means that no change of firmware-level code is required, the code is easily modifiable, and transition from simulation to flight is seamless. Pairing this with powerful tools like Docker [9] simplifies testing and development transitions from vehicle to vehicle. The ROSplane architecture is described more fully in Section III.
ROS 2 has been extensively used for research and commercial use alike[7]. Existing autopilots like PX4 have made significant efforts to support ROS 2 software, and integration with ROS 2 software and PX4 is constantly improving [1]. ROSplane takes ROS 2 support a step further by structuring all autopilot features as ROS 2 nodes, including controllers, estimators, and path planners. This tight integration with ROS 2 significantly improves customizability and modularity.
Another challenge with UAV research is the lack of defined aerodynamic models for many UAVs. This limits the ability of researchers to perform accurate testing in simulation. Further, accurate aerodynamic analysis software is often expensive, which only exacerbates barriers to entry in UAV research. Some recent works have explored using open source aerodynamic modeling and analysis tools for UAV system identification [4], [14]. However, there are few clear documentation and tutorials that assist researchers through the whole process of designing the model, performing stability and control analyses, executing in simulation, and transitioning to hardware. ROSplane 2.0 provides a simulation environment with easily modifiable aerodynamic parameters. Additionally, ROSflight 2.0 provides an aerodynamic modeling pipeline with complete documentation that takes users through this process using two readily-available tools, XFlyer5 [15] and OpenVSP [13].
@@ III-B Autopilot Structure
This architecture gives ROSplane basic waypoint-following functionality out-of-the-box. Due to the design philosophy of ROSplane, the waypoint-following functionality is neither complex, advanced, nor fully-featured. Instead, the structure of ROSplane facilitates easy understanding with clearly delineated responsibilities assigned to each module and lean, well-understood implementations. This makes it easier for users to identify which modules are necessary for their specific applications and which modules should be removed, replaced, or modified, thus increasing modularity and customizability while reducing time invested in deployment.
@@ V-A State Estimation
Previously, ROSplane used a pedagogically-motivated two-stage estimation scheme based on the implementation in [3]. It separated attitude estimation and positional states. This made it easier for students to implement [5]; however, the estimator was less accurate than desired. Additionally, while the implementation in [3] required a measurement of heading (e.g. digital compass or magnetometer), the previous ROSplane estimator did not implement one, resulting in poor wind estimates. In ROSplane 2.0, the estimator has been replaced with a full-state extended Kalman filter (EKF). This full-state EKF has an expanded number of estimated states, including heading, gyroscope biases and lateral velocity. The estimator also takes full advantage of GNSS velocity measurements and measurements from the magnetometer. The updated estimator shares its structure and derivation with the one presented in ROScopter with the addition of two measurement updates [10]. The measurement updates included in ROSplane are the updates for the differential pressure and the zero side-slip angle assumption that drive wind estimation.
The sensor reading $z_{\text{diff}}$ measures the differential pressure from the pitot tube. The measurement model for the differential pressure update $h_{\text{diff}}(x)$, derived from the state $\boldsymbol{x}$, is given by
$h_{\text{diff}}(\boldsymbol{x)}=\frac{1}{2}\rho\big(\boldsymbol{v}-R^{\top}(\boldsymbol{\theta})\boldsymbol{w}\big)^{\top}\big(\boldsymbol{v}-R^{\top}(\boldsymbol{\theta})\boldsymbol{w}\big),$ (1)
where $\boldsymbol{v}$ is the estimated velocities in the body frame, $R(\boldsymbol{\theta})$ is the rotation matrix that rotates values in the body frame into the inertial frame and $\boldsymbol{w}$ is the estimated wind in the inertial frame.
To enforce the nominal dynamics of the aircraft that the side-slip angle $\beta$ is zero, the estimator uses a pseudo-measurement update. This assumes that all velocity along the body y-axis of the airplane is caused by wind. The pseudo-measurement $z_{\beta}$ is taken to be zero and the measurement model $h_{\beta}(\boldsymbol{x})$ is given as,
$h_{\beta}(\boldsymbol{x})=\begin{bmatrix}0&1&0\end{bmatrix}(\boldsymbol{v}-R(\boldsymbol{\theta})^{\top}\boldsymbol{w}).$ (2)
Eq. (2) finds the estimated side-slip velocity while taking away the effects of wind, which should nominally be zero. These measurement updates are incorporated as in [10] and drive the wind estimation portion of the ROSplane estimator.
The EKF also has modularity improvements. Beyond being completely replaceable with the simple output interface, the EKF is designed to easily accommodate new measurement updates, allowing the researcher to easily integrate new sensors. The framework for adding new sensors can be found on the project website.
@@ V-B Control
The controller’s modularity improvements allow for safe integration of new control schemes. This includes utilizing a state machine to ensure that full autonomous control is only taken under safe conditions. ROSplane 2.0 has a simple state machine to allow for the controller to have different control schemes based on the altitude and phase of flight. These are summarized in Figure 3. Integration with these phases of flight and the modular nature of the controller maximize flexibility for the researcher to integrate research by allowing them to take as much or as little control as they need. A researcher can allow ROSplane’s native behavior to dominate in climb then transfer control to a new in development algorithm only when cruise has been achieved.
@@ VI-C Intended Workflow for ROSplane
This section describes the intended workflow for users interested in ROSplane. The first step is for a researcher to develop their application-specific code and integrate it into the ROSplane autonomy stack. Then, the aircraft used in hardware flight tests should be modeled as described in Section IV. If hardware tests will not be performed, then the default aerodynamic model can be used. It is recommended that users then use the ROSflight simulation to test and further refine the application code. The project website has detailed information on setting up the simulation environment. Once the application code works well in simulation, hardware tests can be performed. Because ROSflight and ROSplane run the exact same code in simulation and in hardware [11], these hardware tests can be performed in the exact same manner as the simulation tests were performed.
@@ VII Results
This section presents the results of simulation and hardware experiments demonstrating the basic functionality ROSplane offers to users out-of-the-box. These results also demonstrate that ROSplane closes the simulation-to-real gap for research.
VII-A Simulation
The ROSplane 2.0 full-state estimator was first compared to the previous ROSplane estimator in a simulation environment. An aircraft was flown manually in simulation and the sensor data was recorded and fed through both estimators. Figure 4 compares the errors in estimated state against ground truth. Wind estimates are shown only for the ROSplane 2.0 estimator due to poor performance in the previous version which did not implement a magnetometer. Table III shows RMS error between all estimated states for both estimators. The full-state EKF has significantly less tracking error in attitude, velocity and position, and demonstrates that it provides consistent and reliable estimation to support the flight experiments of researchers. In the companion paper for ROScopter, the improved estimator is shown to perform well compared against the PX4 estimator [10].
Figure 4: Comparison of ROSplane and ROSplane 2.0 estimated state errors during a manually controlled flight in simulation.
@@ VII-B Simulation to Real Flight
The simulation capabilities of ROSplane and the aerodynamic modeling pipeline were tested according to the aircraft integration process outlined in Section IV. The responses of the controller in simulation were then compared to those recorded during a physical flight, given the same trajectory commands and controller configuration.
@@
A model of the RMRC Anaconda UAV shown in Figure 5 was created in both XFLR5 and OpenVSP, and a variety of aerodynamic and stability analyses were performed to obtain the aerodynamic coefficients and stability and control derivatives for the aircraft. XFLR5 and OpenVSP both produced aerodynamic models that were accurate enough for controller tuning, though each tool excelled in different areas. XFLR5’s built-in trimming capability made stability analyses straightforward and yielded more consistent stability derivatives. OpenVSP, by modeling the full aircraft geometry, provided more reliable drag estimates and control derivatives. For both the Xflr5 and OpenVSP models, several cycles of iteration, as outlined in Section IV, were required for the aircraft model to converge on realistic performance. Figure 6 shows the OpenVSP simulation results compared with several circuits from physical flight results. The final performance of the simulation and physical flight was comparable. Note that the same controller gains were used in both flights, meaning that there was no difference in ROSplane’s controllers for this comparison. The Xflr5 results were similar.
@@
These comparisons validate that new UAVs can be modeled and analyzed using open-source software, integrated into the ROSplane simulation, and used for controller tuning, mission planning, and research testing. Iterative testing and improvement of these models provided confidence that real-life system performance can be accurately predicted within a safe simulation environment. This allows for high confidence in the success of research algorithms in real flight tests.
VII-C Hardware
To validate the performance of ROSplane, a simple waypoint mission was flown on the RMRC Anaconda UAV shown in Figure 5. ROSflight firmware was run on an FCU using the configuration 2 hardware [11], which includes an onboard Jetson Orin as the companion computer. In each test, the aircraft was directed to fly to a series of four waypoints in an hourglass formation on a continuous circuit, as shown in Figure 8. RTK GPS data was collected during the flight to validate position estimates, but was not used as input to ROSplane’s estimator. The desired path and ROSplane’s performance is shown in Figure 8, with an estimated 4-6 m/s wind from the south-east direction. The path following algorithm overshoots the turns of the path because the commanded path assumes that the aircraft is only acceleration-limited and can achieve its maximum bank angle instantaneously. The deviations from the commanded altitude in Figure 8 correspond to periods during turn maneuvers, where a drop in altitude is expected due to the loss of lift when rolled.
Figure 9 shows the estimated states and the control setpoints for roll, pitch, course, altitude, and airspeed commands. The aircraft was able to track the given commands effectively. Times of poor response in the pitch loop are coincident with high roll commands, which is expected behavior for pitch-based altitude control. It is important to note that the oscillations seen in the roll response have a frequency on the order of five seconds and are in response to disturbances in the course from the commanded trajectory.
Table I lists the RMS error between ROSplane’s estimated positional states and RTK GPS, showing that ROSplane effectively estimated the vehicle’s position. Figure 7 Ground truth for attitude and airspeed was not available. Altitude error is significantly higher due to both barometric effects of wind speed on the cabin pressure and reduced vertical GPS accuracy. Table II shows RMS error of ROSplane’s estimated state and the commanded trajectory, showing that ROSplane successfully tracked the commanded setpoints. ROSplane is able to achieve consistent and reliably effective results in estimation and waypoint following. These results show that ROSplane has a reliable basic feature set to be an effective platform for research. Researchers can implement desired algorithms easily to evaluate different control, estimation, trajectory planning and following algorithms. Full discussion and analysis of the controller and rationale for its performance can be found in [3].
North (m) East (m) Down (m) Total (m)
RMSE 1.32 1.41 5.64 5.96
TABLE I: Position RMS errors between ROSplane’s estimated state and RTK GPS.
Figure 7: Error between the estimated position state and the RTK GPS data during portions of the flight with RTK GPS lock. RTK GPS was not used as an input to the estimator.
North (m) East (m) Altitude (m) Roll (deg) Pitch (deg) Course (deg) Airspeed (m/s)
RMSE 3.05 3.29 1.19 7.55 3.04 8.51 0.58
TABLE II: RMS errors between the commanded trajectory and the estimated state during four loops of the waypoint mission flown in hardware and shown in Figure 8.
Estimator North (m) East (m) Altitude (m) Vel-X (m/s) Vel-Y (m/s) Vel-Z (m/s) Roll (deg) Pitch (deg) Course (deg)
ROSplane 0.586 0.418 3.974 2.17 0.684 4.820 1.814 5.409 0.702
ROSplane 2.0 0.325 0.258 0.894 0.045 0.089 0.105 0.189 0.074 0.112
TABLE III: RMS errors between estimated and simulated ground truth. The ROSplane 2.0 estimator significantly reduces RMS errors in estimated state.
Figure 8: ROSplane estimated position alongside commanded path. Moderate winds from the south-east were present. ROSplane is able to successfully follow the waypoint mission.
Figure 9: The controller response to given flight path input in the attitude angles, airspeed, and altitude. ROSplane effectively tracks the commanded setpoints.
@@ VIII Conclusion
ROSplane’s structure, capabilities, and modularity provide researchers with more flexibility and lower the barriers to entry to perform UAV autonomy research. The improved code structure and architecture provide greater customizability and opportunities for integrating external hardware and code. ROSplane 2.0 architecture allows researchers to quickly and safely tune flight controllers and test algorithms for new fixed-wing UAV research in simulation to prepare for physical flight tests. This functionality of ROSplane 2.0 was demonstrated by modeling a UAV aircraft using open-source aircraft modeling and analysis software, integrating this aircraft model into the ROSplane simulation, and validating that the closed-loop controller performance of the simulation and physical flight are comparable. This enables researchers to perform accurate controller tuning and algorithm testing in simulation in the actual flight control software before flight, thereby reducing risk and iteration time. The universal applicability of the ROS 2 structure, paired with proven functionality and an effective simulation platform, makes ROSplane a powerful tool for researchers valuing customizability and direct control in cutting-edge autonomous research applications.